# Exploration and Incentivizing Participation in Clinical Trials

### v8

@@ 3.1 Benchmark
The benchmark is finite as long as the best arm strictly improves over the outside option. [^17] Failing that, we have a degenerate case when the patients believe that the trial is completely useless. To express this point, let $a^{*}(\psi)\in\operatornamewithlimits{argmax}_{a\in[n]}U_{\psi}(a)$ be the best arm: a utility-maximizing action for a given state $\psi$, ties broken arbitrarily. Its Bayesian-expected reward is $U^{*}_{\mathcal{P}}:=\operatornamewithlimits{\mathbb{E}}_{\psi\sim\mathcal{P}}\left[\,U_{\psi}\left(\,a^{*}(\psi)\,\right)\,\right]$.
@@ 3.4 Lower bound
Thus, Theorem 3.7 allows per-round policies that change over time and can themselves be jointly chosen at random. In rounds $t\not\in S$ the policies can depend on all past observations. Importantly, this includes “multi-stage” mechanisms that partition the time into a constant number of contiguous “stages” and are non-data-adaptive within each (possibly with a data-adaptive warm-up stage of duration, say, $<T/2$). Most medical trials in practice fit this model, see Footnote 7.
@@ 4 Heterogeneous agents
We now consider the general case of heterogeneous agents. Agent heterogeneity is two-fold: different agent types may have different outcome distributions for the same arm, and different subjective (and not directly observable) utilities for the same outcome. We model this as follows: public types determine outcome distributions and the private types determine subjective utilities. The intuition is that beliefs about objective outcomes are driven by the patient’s objective medical history, which is typically well-documented and available to the clinical trial. However, patients’ preferences over outcome distributions can be highly subjective, e.g., in evaluating the disutility of particular medical complications. These subjective preferences are not known to the mechanism. In our model, each agent reports its private type upon arrival, and the mechanism must incentivize truthful reporting (in addition to the BIR property).
@@ 4.3 Our results
Warm-up stage. Warm-up data collection is treated separately, as in Section 3. It can be done either endogenously (via a BIR, BIC mechanism), or exogenously (e.g., by paying the volunteers). Either way, the sufficient amount of warm-up data depends only on the prior and the utility structure, whereas it bootstraps our mechanism to run for an arbitrarily large time horizon $T$.
Endogenous data collection has been studied in prior work on incentivized exploration, for both public and private agent types. In particular, for public types one can collect samples for all type-arm pairs, under some assumptions (Mansour et al.,, 2020). For arbitrary types, one can, in some sense, explore all type-arm pairs that can possibly be explored (Immorlica et al.,, 2019). Recall that a sufficient amount of warm-up data needs to be guaranteed only w.r.t. agents’ beliefs, in particular, only for stochastic adversaries.
@@
$\geq U_{\mathtt{mech}}(\psi^{*},\theta)-\Pr\left[\,\psi_{\mathtt{MLE}}\neq\psi^{*}\,\right]\geq U_{\mathtt{out}}(\theta).$
@@
$\geq$ $\,U_{\mathtt{mech}}(\psi_{\mathtt{MLE}},\theta,\theta^{\prime})+\frac{1}{2}\cdot\eta-2\Pr\left[\,\psi_{\mathtt{MLE}}\neq\psi^{*}\,\right]\geq U_{\mathtt{mech}}(\psi_{\mathtt{MLE}},\theta,\theta^{\prime}).$
@@
$\leq\frac{1}{(T-T_{0})^{2}}\;\max_{a\in[n]}\max_{\psi}\sum_{\theta}\sum_{t\in[T_{0},T]}\frac{{\bf 1}_{\left\{\,\theta_{t}=\theta\,\right\}}}{\frac{1}{2}\cdot\sigma^{\mathtt{opt}}_{a}(\psi,\theta)}$
@@ 5 Heterogeneous agents with estimated type frequencies
In this section, we consider heterogeneous agents from a different perspective, whereby we fix the distribution $F_{\mathtt{type}}$ of type frequencies. The benchmark $\mathtt{bench}(\mathcal{P},\Theta)$ in Section 4 takes a $\max$ over all types (see Eq. (4.8)), and may therefore be skewed towards “bad” types – ones with large benchmark value – that are difficult for the mechanism to handle. To mitigate this issue when the bad types are rare, we define an alternative benchmark parameterized by $F_{\mathtt{type}}$, denoted $\mathtt{bench}(\mathcal{P},F_{\mathtt{type}})$, which replaces the $\max$ over all types with an expectation over $F_{\mathtt{type}}$. Our mechanism attains guarantees with respect to this benchmark, when initialized with estimated type frequencies $\widetilde{F}_{\mathtt{type}}$ that are not too far from $F_{\mathtt{type}}$. Such estimates can be constructed exogenously via surveys, medical records, or past clinical trials). When $\widetilde{F}_{\mathtt{type}}=F_{\mathtt{type}}$, this guarantee is optimal, much like in Theorem 3.7, in the worst case over all adversaries with a given $F_{\mathtt{type}}$.
With this approach, the influence of rare-but-difficult agent types can now be understood solely from analyzing the benchmark $\mathtt{bench}(\mathcal{P},F_{\mathtt{type}})$. The issue is quite intricate, however, because “difficulty” of a type for our benchmark is not well-defined in isolation from the other types. Instead, it is a property of the entire distribution over types. In Section 5.3, we consider a simple scenario with two distributions over types, “good” and “bad”, and mixtures thereof. We analyze the benchmark and show that the influence of the “bad” distribution over types is indeed mitigated.
@@ 5.1 Results
The realized type frequencies are defined over the main stage, in a natural way:
[…]
We posit that the mechanism is initialized with estimated type frequencies: a distribution over types $\widetilde{F}_{\mathtt{type}}$ that estimates $F_{\mathtt{type}}(\cdot\mid\mathtt{adv})$ for the (actual) adversary $\mathtt{adv}$. Such estimates can be constructed exogenously via surveys, medical records, past clinical trials, etc.
Remark 5.1.
We only need to approximate the empirical frequency of types. That is, the type sequence is still chosen by the adversary. Our guarantees do not require the agent types to be drawn i.i.d. in each round.
The new benchmark replaces the $\max$ over all types in (4.7) with an expectation:
Remark 5.2.
The new benchmark (5.2) specializes to Eq. (3.1) for homogeneous agents, just like the worst-case benchmark $\mathtt{bench}(\mathcal{P},\Theta)$. However, (5.2) does not appear to admit a clean alternative characterization in terms of the per-type benchmarks. In particular, we do not have $\mathtt{bench}(\mathcal{P},F_{\mathtt{type}})=\operatornamewithlimits{\mathbb{E}}_{\theta\sim F_{\mathtt{type}}}\mathtt{bench}(\mathcal{P},\{\theta\}).$
It is easy to see that $\mathtt{bench}(\mathcal{P},F_{\mathtt{type}})\leq\mathtt{bench}(\mathcal{P},\Theta)$. (Indeed, the only difference between their resp. definitions, (5.2) and (4.7), is that the max over types is replaced with an expectation.) $\mathtt{bench}(\mathcal{P},F_{\mathtt{type}})$. In Appendix A we provide a simple numerical example when $\mathtt{bench}(\mathcal{P},F_{\mathtt{type}})$ provides a substantial improvement.
In the subsequent developments, we reuse much of the machinery developed in Section 4. For starters, we re-use the non-degeneracy condition (4.9), with the same proof as in Claim 4.7.
@@
We reuse mechanism $\mathtt{TwoStageMech}(T_{0},\sigma^{\mathtt{opt}},a^{*}())$, where $a^{*}()$ is the best-arm policy and $\sigma^{\mathtt{opt}}$ is the benchmark-optimizing policy. The latter is defined relative to the estimated type distribution, $\sigma^{\mathtt{opt}}=\sigma^{\mathtt{opt}}(\mathcal{P},\widetilde{F}_{\mathtt{type}})$. The number of warm-up samples is given by the same Eq. (4.12).
@@
$=\mathtt{bench}(\mathcal{P},F_{\mathtt{type}})+C_{\mathcal{P}}(F_{\mathtt{type}})\cdot\lVert\,\widetilde{F}_{\mathtt{type}}-F_{\mathtt{type}}\,\rVert_{1}.$
@@ 5.3 Vanishing impact of bad types
Let us analyze benchmark $\mathtt{bench}(\mathcal{P},F_{\mathtt{type}})$ to clarify a key advantage thereof: the ability to limit the influence of agent types that are rare but difficult to incentivize. This point is somewhat intricate to formulate, because $\mathtt{bench}(\mathcal{P},F_{\mathtt{type}})$ is not readily expressible in terms of the per-type benchmarks (see Remark 5.2).
We focus on a relatively simple scenario with two distributions over types, $F_{\mathtt{good}}$ and $F_{\mathtt{bad}}$. We posit that $\mathtt{bench}(\mathcal{P},F_{\mathtt{bad}})$ is “large” and $\mathtt{bench}(\mathcal{P},F_{\mathtt{bad}})$ is “small”. The base case is that $F_{\mathtt{good}},F_{\mathtt{bad}}$ are each supported on a single type: resp., a “good type” and a “bad type”. More generally, $F_{\mathtt{good}}$ (resp., $F_{\mathtt{bad}}$) can be supported on a subset of “good types” (resp., “bad types”); further, we allow their support sets to overlap on some “medium types”.
@@ 2. Option 2:
Thus any type $\theta\in{\rm supp}(F_{\mathtt{good}})$ will prefer option 1 to option 2. Moreover, given option $1$, type $\theta$ will not have incentives to deviate his report since both $\sigma^{\mathtt{good}}$ and $\sigma^{\eta}$ are BIC for the type space ${\rm supp}(F_{\mathtt{good}})$, and deviating the report to types in ${\rm supp}(F_{\mathtt{bad}})\backslash{\rm supp}(F_{\mathtt{good}})$ will lead to outside option value, which is not beneficial for the patient.
@@ 6 Conclusions and open questions
We consider three model variants: homogenous agents, heterogenous agents with no ancillary information, and heterogeneous agents with estimated type frequencies. The third variant mitigates the influence of rare-but-difficult agent types. For each model variant, we identify an abstract “comparator benchmark”, design a simple mechanism which attains this benchmark, and prove that no mechanism can achieve better statistical performance. Our results focus on mechanisms that do not adapt to data over the main stage, as defined in Section 2.4 and motivated by practical considerations.
Our work sets the stage for further investigation. First, one could ask concrete questions that go slightly beyond our modelling approach. Can one match our guarantees for heterogeneous agents if the type space is not fully known, or the type frequencies are known very imprecisely? As the duration of the warm-up stage could be prohibitively large in practice, particularly in the presence of “difficult types”, can one reduce it and/or use some of its data for the counterfactual estimation? Second, one could ask a more “conceptual” question: what if the patients cannot fully understand or quantify their risk preferences, i.e., their own private types? What would be more realistic ways to elicit their preferences, and how to incorporate that into incentivized medical trials?
@@ Appendix A Simple numerical examples
Homogeneous agents. There are $n=2$ treatments (henceforth, “arms”), the “old” and the “new”. As per our model, agents believe each arm gives the same utility distribution for all patients, and the state (which we identify with a mapping from arms to utility distributions) is initially drawn from some prior $\mathcal{P}$ (shared by all agents). Specifically, the old arm gives mean utility 1 with probability $\nicefrac{{2}}{{3}}$ and utility 0 with probability $\nicefrac{{1}}{{3}}$. The new arm gives utility 2 and -1 with probability $\nicefrac{{1}}{{2}}$ each. The prior is independent across arms. Given the mean utilities (and the fact that the realized utilities lie in the $[0,1]$ interval), the noise shape can be arbitrary for this example. [^25] Thus, in the formalism of our model, we have four possible states, with mean utility pairs $(1,2)$, $(1,-1)$, $(0,2)$ and $(0,-1)$, and respective prior probabilities $\nicefrac{{1}}{{3}}$, $\nicefrac{{1}}{{3}}$, $\nicefrac{{1}}{{6}}$, $\nicefrac{{1}}{{6}}$. Agents’ alternative to participation is to use the old arm, whose prior-mean utility is $\nicefrac{{2}}{{3}}$; so, the “outside option” is $U_{\mathtt{out}}=\nicefrac{{2}}{{3}}$. Note that it is ex-ante preferred to the new arm (whose prior-mean utility is only $\nicefrac{{1}}{{2}}$), so the agents are incentivized against participation if the arm probabilities are data-independent.
Our participation-incentivizing RCT design from Section 3 works out as follows. Leveraging Lemma 3.2, consider $\epsilon^{*}$, the largest $\epsilon$ such that $(\epsilon,\psi^{*})\text{-$\mathtt{greedy}$}$ policy is BIR. We claim $\epsilon^{*}=\nicefrac{{8}}{{9}}$. Indeed, note that $\epsilon^{*}$ is is the unique $\epsilon$ such that the Bayesian-expected utility of $(\epsilon,\psi^{*})\text{-$\mathtt{greedy}$}$, call it $U(\epsilon)$, equals the outside option $U_{\mathtt{out}}$. And one can check that $U(\nicefrac{{8}}{{9}})=\nicefrac{{2}}{{3}}=U_{\mathtt{out}}$. (One way to do that is to sum over the four possible states, as per Footnote 25, and for each state calculate the expected utility of the policy; we omit the easy details.)
@@
Now let us calculate $N_{\mathcal{P}}$, the number of samples required for the warm-up stage. The prior-expected utility of the best arm is $U_{\mathcal{P}}(a^{*}())=\nicefrac{{4}}{{3}}$ (the calculation is similar to that for $U(\epsilon^{*})$.) So, the “prior gap” (see Eq. (3.2)) is $\mathtt{gap}(\mathcal{P})=\nicefrac{{4}}{{3}}-\nicefrac{{5}}{{12}}=\nicefrac{{11}}{{12}}$. Therefore, according to Eq. 3.3, we have $\alpha=\nicefrac{{22}}{{27}}$ and $N_{P}<144$. In other words, a warm-up stage with $144$ independent samples of each arm suffices to guarantee incentive-compatibility for the main stage.

### v9

@@ 3.1 Benchmark
The benchmark is finite as long as the best arm strictly improves over the outside option. [^20] Failing that, we have a degenerate case when the agents believe that the trial is completely useless. To express this point, let $a^{*}(\psi)\in\operatornamewithlimits{argmax}_{a\in[n]}U_{\psi}(a)$ be the best arm: a utility-maximizing action for a given state $\psi$, ties broken arbitrarily. Its Bayesian-expected reward is $U^{*}_{\mathcal{P}}:=\operatornamewithlimits{\mathbb{E}}_{\psi\sim\mathcal{P}}\left[\,U_{\psi}\left(\,a^{*}(\psi)\,\right)\,\right]$.
@@ 3.4 Lower bound
Thus, Theorem 3.7 allows per-round policies that change over time and can themselves be jointly chosen at random. In rounds $t\not\in S$ the policies can depend on all past observations. Importantly, this includes “multi-stage” mechanisms that partition the time into a constant number of contiguous “stages” and are non-data-adaptive within each (possibly with a data-adaptive warm-up stage of duration, say, $<T/2$). Most medical trials in practice fit this model, see Footnote 10.
@@ 4 Heterogeneous agents
We now consider the general case of heterogeneous agents. Agent heterogeneity is two-fold: different agent types may have different outcome distributions for the same arm (according to the agents’ beliefs), and different subjective utilities for the same outcome (which are not directly observable). We model this as follows: public types determine outcome distributions and the private types determine subjective utilities. The intuition is that beliefs about objective outcomes are driven by the patient’s objective medical history, which is typically well-documented and available to the clinical trial, whereas patients’ (dis)utilities over medical outcomes can be highly subjective. For online platforms, we posit that users’ beliefs about their own objective outcomes are driven by their attributes that are visible to the platform (via user profiles, cookies, etc.), whereas the respective utilities (such as value-per-click) are subjective and unobservable. Each agent reports its private type upon arrival, and the mechanism must incentivize truthful reporting.
@@ 4.3 Our results
Warm-up stage is treated separately, as in Section 3, either endogenously (via a BIR, BIC mechanism), or exogenously (e.g., via paid volunteers). The sufficient amount of warm-up data depends only on the agent beliefs and the utility structure, whereas it bootstraps our mechanism to run for an arbitrarily large time horizon $T$. Endogenous data collection has been studied in prior work on incentivized exploration, for both public and private types: in some sense, one can explore all type-arm pairs that can possibly be explored (Mansour et al.,, 2020; Immorlica et al.,, 2019).
@@
$\geq U_{\mathtt{mech}}(\psi^{*},\theta)-\Pr\left[\,\psi_{\mathtt{MLE}}\neq\psi^{*}\,\right]$
$\geq U_{\mathtt{out}}(\theta).$
@@
$\geq\,U_{\mathtt{mech}}(\psi_{\mathtt{MLE}},\theta,\theta^{\prime})+\frac{1}{2}\cdot\eta-2\Pr\left[\,\psi_{\mathtt{MLE}}\neq\psi^{*}\,\right]$
$\geq U_{\mathtt{mech}}(\psi_{\mathtt{MLE}},\theta,\theta^{\prime}).$
@@
$\leq\frac{1}{(T-T_{0})^{2}}\;\max_{a\in[n]}\max_{\psi}\sum_{\theta}\sum_{t\in[T_{0},T]}\frac{{\bf 1}_{\left\{\,\theta_{t}=\theta\,\right\}}}{\frac{1}{2}\cdot\widehat{\sigma}^{\mathtt{opt}}_{a}(\psi,\theta)}$
@@ 5 Heterogeneous agents with estimated type frequencies
This section focuses on type frequencies $F_{\mathtt{type}}$ in the main stage:
[…]
The mechanism is initialized with estimated type frequencies $\widetilde{F}_{\mathtt{type}}$: a distribution over types that estimates $F_{\mathtt{type}}(\cdot\mid\mathtt{adv})$ for the (actual) adversary $\mathtt{adv}$. Such estimates can be constructed exogenously by a clinical trial (via surveys, medical history, or data from the past clinical trial) or an online platform (via user profiles and their interaction history).
The goal is to mitigate the influence of rare-but-difficult types. The benchmark $\mathtt{bench}(\mathcal{P},\Theta)$ in Section 4 (see Eq. (4.8)) is determined by the “bad” types (ones with large benchmark value) that are difficult for the mechanism to handle. To mitigate this when the bad types are rare, we define an alternative benchmark which replaces the $\max$ over all types with an expectation over $F_{\mathtt{type}}$:
We provide upper and lower bounds relative to this benchmark. Our mechanism attains a strong guarantee when the estimates $\widetilde{F}_{\mathtt{type}}$ are not too far from $F_{\mathtt{type}}$. When $\widetilde{F}_{\mathtt{type}}=F_{\mathtt{type}}$, this guarantee are optimal, much like in Theorem 3.7, in the worst case over all adversaries with a given $F_{\mathtt{type}}$.
The influence of rare-but-difficult agent types can now be understood solely from analyzing the benchmark $\mathtt{bench}(\mathcal{P},F_{\mathtt{type}})$. It is easy to see that $\mathtt{bench}(\mathcal{P},F_{\mathtt{type}})\leq\mathtt{bench}(\mathcal{P},\Theta)$. In Appendix A we provide a simple numerical example when $\mathtt{bench}(\mathcal{P},F_{\mathtt{type}})$ provides a substantial improvement. However, we do not have $\mathtt{bench}(\mathcal{P},F_{\mathtt{type}})=\operatornamewithlimits{\mathbb{E}}_{\theta\sim F_{\mathtt{type}}}\mathtt{bench}(\mathcal{P},\{\theta\})$, and more generally $\mathtt{bench}(\mathcal{P},F_{\mathtt{type}})$ does not appear to admit a clean characterization in terms of the per-type benchmarks. Instead, the influence of a particular type on the benchmark depends on the entire distribution over types. The entire distribution could be “bad”, not just the individual types.
Thus, analyzing the influence of “bad” types/distributions beyond numerical examples is quite subtle. In Section 5.3, we consider a scenario with two type distributions, “good” and “bad”, and their mixtures. We analyze the benchmark and show that the influence of the “bad” type distribution on the mixture is indeed mitigated, in some asymptotic sense.
@@ 5.1 Results
While we add some new ideas, we reuse much of the machinery developed in Section 4. For starters, we re-use the non-degeneracy condition (4.9), with the same proof as in Claim 4.7.
@@
We reuse mechanism $\mathtt{TwoStageMech}(T_{0},\sigma^{\mathtt{opt}},a^{*}())$ from Section 4 where $a^{*}()$ is the best-arm policy and $\sigma^{\mathtt{opt}}=\sigma^{\mathtt{opt}}(\mathcal{P},\widetilde{F}_{\mathtt{type}})$ is the benchmark-optimizing policy relative to the estimated type frequencies. The number of warm-up samples is given by the same Eq. (4.12).
@@
$=\mathtt{bench}(\mathcal{P},F_{\mathtt{type}})+C_{\mathcal{P}}(F_{\mathtt{type}})\cdot\lVert\,\widetilde{F}_{\mathtt{type}}-F_{\mathtt{type}}\,\rVert_{1}.\qed$
@@ 5.3 Vanishing impact of bad types
Let us analyze benchmark $\mathtt{bench}(\mathcal{P},F_{\mathtt{type}})$ to clarify its key advantage: the ability to limit the influence of “bad” types (or “bad” distributions over types). This is an intricate point, as discussed early in this section.
We focus on a simple scenario with two distributions over types, $F_{\mathtt{good}}$ and $F_{\mathtt{bad}}$. We posit that $\mathtt{bench}(\mathcal{P},F_{\mathtt{bad}})$ is “large” and $\mathtt{bench}(\mathcal{P},F_{\mathtt{bad}})$ is “small”. The base case is that $F_{\mathtt{good}},F_{\mathtt{bad}}$ are each supported on a single type: resp., a “good type” and a “bad type”. More generally, $F_{\mathtt{good}}$ (resp., $F_{\mathtt{bad}}$) can be supported on a subset of “good types” (resp., “bad types”); further, we allow their support sets to overlap on some “medium types”.
@@ 2. Option 2:
Thus any type $\theta\in{\rm supp}(F_{\mathtt{good}})$ will prefer option 1 to option 2. Moreover, given option $1$, type $\theta$ agent will not have incentives to deviate his report since both $\sigma^{\mathtt{good}}$ and $\sigma^{\eta}$ are BIC for the type space ${\rm supp}(F_{\mathtt{good}})$, and deviating the report to types in ${\rm supp}(F_{\mathtt{bad}})\backslash{\rm supp}(F_{\mathtt{good}})$ will lead to this agent receiving value $U_{\mathtt{out}}$, which is not beneficial for the agent.
@@ 6 Conclusions and open questions
We consider three model variants: homogeneous agents, heterogeneous agents with no ancillary information, and heterogeneous agents with estimated type frequencies. For each model variant, we identify an abstract “comparator benchmark”, design a simple mechanism which attains this benchmark, and prove that no mechanism can achieve better statistical performance. Our mechanisms have a desirable property of not adapting to the data collected after the “warm-up stage”; however, our negative results apply to a much wider class of mechanisms.
Our work sets the stage for further investigation. First, while we mitigate the influence of rare-but-difficult agent types on the main stage (via estimated type frequencies), can one do it for the warm-up stage, too? That is, reduce its duration and/or reuse its data for the counterfactual estimates. Second, can one match the result for known type frequencies if the type frequencies are instead learned on the fly? Third, how to think about incentivized RCTs if the agents cannot fully quantify their own risk preferences, and therefore cannot fully report them?
@@ Appendix A Simple numerical examples
Homogeneous agents. There are $n=2$ treatments (henceforth, “arms”), the “old” and the “new”. As per our model, agents believe each arm gives the same utility distribution for all agents, and the state (which we identify with a mapping from arms to utility distributions) is initially drawn from some prior $\mathcal{P}$ (shared by all agents). Specifically, the old arm gives mean utility 1 with probability $\nicefrac{{2}}{{3}}$ and utility 0 with probability $\nicefrac{{1}}{{3}}$. The new arm gives utility 2 and -1 with probability $\nicefrac{{1}}{{2}}$ each. The prior is independent across arms. Given the mean utilities (and the fact that the realized utilities lie in the $[0,1]$ interval), the noise shape can be arbitrary for this example. [^28] Thus, in the formalism of our model, we have four possible states, with mean utility pairs $(1,2)$, $(1,-1)$, $(0,2)$ and $(0,-1)$, and respective prior probabilities $\nicefrac{{1}}{{3}}$, $\nicefrac{{1}}{{3}}$, $\nicefrac{{1}}{{6}}$, $\nicefrac{{1}}{{6}}$. Agents’ alternative to participation is to use the old arm, whose prior-mean utility is $\nicefrac{{2}}{{3}}$; so, the “outside option” is $U_{\mathtt{out}}=\nicefrac{{2}}{{3}}$. Note that it is ex-ante preferred to the new arm (whose prior-mean utility is only $\nicefrac{{1}}{{2}}$), so the agents are incentivized against participation if the arm probabilities are data-independent.
Our participation-incentivizing RCT design from Section 3 works out as follows. Leveraging Lemma 3.2, consider $\epsilon^{*}$, the largest $\epsilon$ such that $(\epsilon,\psi^{*})\text{-$\mathtt{greedy}$}$ policy is BIR. We claim $\epsilon^{*}=\nicefrac{{8}}{{9}}$. Indeed, note that $\epsilon^{*}$ is is the unique $\epsilon$ such that the Bayesian-expected utility of $(\epsilon,\psi^{*})\text{-$\mathtt{greedy}$}$, call it $U(\epsilon)$, equals the outside option $U_{\mathtt{out}}$. And one can check that $U(\nicefrac{{8}}{{9}})=\nicefrac{{2}}{{3}}=U_{\mathtt{out}}$. (One way to do that is to sum over the four possible states, as per Footnote 28, and for each state calculate the expected utility of the policy; we omit the easy details.)
@@
Now let us calculate $N_{\mathcal{P}}$, the number of samples required for the warm-up stage. The prior-expected utility of the best arm is $U_{\mathcal{P}}(a^{*}())=\nicefrac{{4}}{{3}}$ (the calculation is similar to that for $U(\epsilon^{*})$.) So, the “prior gap” (see Eq. (3.2)) is $\mathtt{gap}(\mathcal{P})=\nicefrac{{4}}{{3}}-\nicefrac{{5}}{{12}}=\nicefrac{{11}}{{12}}$. Therefore, according to Eq. 3.3, we have $\alpha=\nicefrac{{22}}{{27}}$ and $N_{P}<144$. In other words, a warm-up stage with $144$ independent samples of each arm suffices to guarantee incentive-compatibility for the main stage. Recall that these samples may also be obtained exogenously, e.g., via paid volunteers.